OpenStudy (anonymous):

How do I solve this a squared + 40 squared = 96 squared

OpenStudy (anonymous):

\[a^2+40^2=96^2 \]\[a^2=96^2-40^2 \]\[a=\sqrt{7616} \]\[a=8 \sqrt{119} \]